Find the slope, x-intercept, and y-intercept for the line 4x-8y-5=0
Answer by fractalier(6550)

The x-intercept is found by setting y equal to 0 and solving for x...
so 4x - 5 = 0 and 4x = 5 and x = 5/4
The y-intercept is found by setting x equal to 0 and solving for y...
so -8y - 5 = 0 and -8y = 5 and y = -5/8
Since it is in standard form Ax + By + C = 0, the slope is -A/B = -4/-8 = 1/2